OpenStudy (anonymous):

Factor completely: 14bx2 - 7x3 - 4b + 2x

OpenStudy (anonymous):

So we need to factor the like terms out. 14bx^2-7x^3-4b+2x =14x^2(2b-x)-2(2b-x) =(2b-x)(14x^2-2) =2(2b-x)(7x^2-1) This is called factorizing by grouping. Let me know if any step is unclear.